OpenStudy (anonymous):

DOES ANYONE KNOW HOW TO SIMPLIFY THESE>???? #1)tan x(cos x – csc x) #2)(sin x + cos x)(sec x+csc x)

OpenStudy (anonymous):

great

OpenStudy (eyust707):

first you gotta know some of the identities....

OpenStudy (eyust707):

\[\tan x = { \sin x \over \cos x}\]

OpenStudy (eyust707):

\[\sec x = { 1 \over \cos x}\]

OpenStudy (eyust707):

\[\csc x = { 1 \over \sin x}\]

OpenStudy (eyust707):

Lets do number one

OpenStudy (anonymous):

alright

OpenStudy (eyust707):

change the tan and the csc to the things i just taught you

OpenStudy (eyust707):

I get: \[{\sin x \over \cos x} (\cos x - {1 \over \sin x})\]

OpenStudy (eyust707):

then distribute in

OpenStudy (anonymous):

thats what i got

OpenStudy (eyust707):

\[\sin x - {1 \over \sin x}\]

OpenStudy (anonymous):

right

OpenStudy (eyust707):

I prefer that form... your teacher may like csc instead

OpenStudy (eyust707):

either way you are finished...

OpenStudy (anonymous):

It's - 1/cos = sec

OpenStudy (anonymous):

what about the second one??

OpenStudy (eyust707):

What I posted was correct

OpenStudy (anonymous):

ok thats what i thought

OpenStudy (eyust707):

For the second one lets leave in sec x and cscx for now and treat it just like algebra

OpenStudy (anonymous):

Sorry to interrupt. You made a mistake multiplying sin/cos times 1/sin in the first problem. It would be 1/cos

OpenStudy (eyust707):

ahh yes you are correct! sorry... \[\sin x - {1 \over \cos x}\]

OpenStudy (anonymous):

I'll leave you to it. You're doing a great job.

OpenStudy (anonymous):

@eyust707 small correction to #1 the solution is \[\sin x - \sec x\] just recheck

OpenStudy (anonymous):

ummmm ok now everyone is getting me confused :(

OpenStudy (eyust707):

=P I made a little mistake on this one: \[{\sin x \over \cos x} (\cos x - {1 \over \sin x})\]

OpenStudy (eyust707):

Notice when you distribute the \(\sin x \over \cos x \) to both terms you end up with \[\sin x - {1 \over \cos x}\] not the other expression i posted previously

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

okay so thats #1

OpenStudy (eyust707):

#2 you just follow the rules of algebra

OpenStudy (anonymous):

alright good :)

OpenStudy (eyust707):

Pretend A B C D are the trig functions of that problem